# Deep learning interpretability analysis for carbon star identification in DR3

### v1

@@ Abstract
Context. A large fraction of Asymptotic Giant Branch (AGB) stars develop carbon-rich atmospheres during their evolution. Based on their color and luminosity, these carbon stars can be easily distinguished from many other kinds of stars. However, a large number of G, K, and M giants are also distributed in the same region as carbon stars on the HR diagram. Their spectra have differences, especially in the prominent CN molecular bands.
Aims. We aim to distinguish carbon stars from other kinds of stars using βs XP spectra, while providing attribution explanations of key features necessary for identification, and even discovering additional new spectral key features.
Methods. We proposed a classification model named βGaiaNetβ, an improved one-dimensional convolutional neural network specifically designed for handling βs XP spectra. We utilized the SHAP interpretability model to calculate the SHAP value for each feature point in a spectrum, enabling us to explain the output of the βGaiaNetβ model and provide further meaningful analysis.
Results. Compared to four traditional machine-learning methods, the βGaiaNetβ model exhibits an average classification accuracy improvement of approximately 0.3% on the validation set, with the highest accuracy even reaching 100%. Utilizing the SHAP algorithm, we present a clear spectroscopic heatmap highlighting molecular band absorption features primarily distributed mainly around CN and CN, and summarize five crucial feature regions for carbon star identification. Upon applying the trained classification model to the CSTAR sample with βxp_sampled_meanβ spectra, we obtained 451 new candidate carbon stars as a by-product.
@@
Carbon stars, first recognized by Secchi (1869), exhibit an inversion of the C/O ratio (C/O ΒΏ 1). Compared to other stars, carbon stars possess carbon-enriched atmospheres . The enrichment can originate from mass transfer in binary systems or the activation of the third dredge-up process during their AGB evolution phase (Gaia Collaboration etΒ al. 2023). The majority of carbon stars in the galaxy are binary systems wherein the host star inherited atmospheric carbon through mass transfer from its AGB companion, which is now a white dwarf (Ji etΒ al. 2016). The carbon enrichment of the cool and luminous carbon stars (C-N type) is due to the pollution by nuclear He fusion products from the inner to the outer layers in their AGB phase, which are called intrinsic carbon stars (Ji etΒ al. 2016). This process transports carbon from the interior region to the stellar surface, leaving the outer layers rich in carbon-containing molecules and dust particles. Thus, their spectra show C and CN molecular bands stronger than usual in stars cooler than 3,800 K (i.e. ) (Gaia Collaboration etΒ al. 2023).
@@
Carbon stars exhibit strong absorption characteristics due to molecular bands of carbon-containing compounds such as C, CN, CH, and various hydrocarbons. The presence and intensity of these spectral features provide valuable insights into the atmospheric conditions and chemical processes within these stars. Carbon stars are also important contributors to the interstellar medium and serve as good reference cases for studying the physical processes affecting the end of the life of low-mass stars (Gaia Collaboration etΒ al. 2023).
@@
Traditionally, identifying and classifying carbon stars relied on parametric measurements and manual checking of their spectra. Ji etΒ al. (2016) identified 894 carbon stars from LAMOST DR2 by measuring multiple line indices from the stellar spectra. Abia etΒ al. (2022) reported the identification of 2,660 new carbon star candidates through 2MASS photometry, astrometry, and their location in the β2MASS diagram. Li etΒ al. (2023) distinguished carbon stars from M-type giants by selecting the criterion of spectral indices satisfying []. Gaia Collaboration etΒ al. (2023) screened carbon stars by measuring the carbon molecular band head strength. Lebzelter etΒ al. (2023) classified 546,468 carbon star candidates using a method guided by narrow-band photometry (Palmer & Wing 1982). Their SOS module examines this feature in an automated way by computing the pseudo-wavelength difference between the two highest peaks in each RP spectrum of a given source, taking the median value of the results, and storing it in the parameter of median_delta_wl_rp. They assumed all stars with median_delta_wl_rp to be C rich (i.e., isCstar = TRUE). Li etΒ al. (2024) identified 3546 carbon stars through line index and near-infrared colorβcolor diagrams. Through visual inspection of these carbon star spectra, they further subclassify them into CβH, CβR, CβN, and Ba stars.
@@
To date, many machine-learning methods have been used to identify carbon stars based on their spectra. Si etΒ al. (2014) applied the label propagation algorithm to search for 260 new carbon stars from SDSS (Sloan Digital Sky Survey) DR8. Si etΒ al. (2015) applied the efficient manifold ranking algorithm to search for 183 carbon stars from the LAMOST pilot survey. Li etΒ al. (2018) identified 2,651 carbon stars in the spectra of more than seven million stars using an efficient machine-learning algorithm for LAMOST (Large Sky Area Multi-Object Fiber Spectroscopic Telescope) DR4. Sanders & Matsunaga (2023) investigated the use of unsupervised learning algorithms to classify the chemistry of long-period variables (Lebzelter etΒ al. 2023) from DR3βs BP/RP spectra (also called XP spectra; Carrasco etΒ al. 2021) into O-rich and C-rich groups. They also employed a supervised approach to separate O-rich and C-rich sources using broadband optical and infrared photometry. In all, they tagged a total of 23,737 C-rich classifications based on the BP/RP spectra and identified a small population of C-rich stars in the Galactic bar-bulge region.
@@
In recent years, the application of interpretable analysis techniques in astronomy has shown great potential. Qin etΒ al. (2019) used a random forest (RF) algorithm deriving rank features, then picked out 15,269 Am candidates from the early-type stars of LAMOST DR5. He etΒ al. (2022) successfully distinguished between RGB and RC stars using the XGBoost algorithm (Chen & Guestrin 2016), and used the SHAP interpretable model (Lundberg & Lee 2017) to obtain the top features that the XGBoost selected. Shang etΒ al. (2022) presented a catalog of 6,917 CP1 and 1,652 CP2 new candidate sources using XGBoost followed by the visual investigation, and listed the spectral features for separating CP1 from CP2 using SHAP.
@@
Before this work, the method of searching for carbon stars based on spectra was mainly to obtain the matching giant spectra by random forest classifier (Breiman 2001), and then to screen out carbon stars with strong CN molecular bands by measuring the molecular band head strength (Gaia Collaboration etΒ al. 2023), actually most of which belong to the N-type. Although this method can easily pick out the carbon stars with obvious spectral features, it also introduces pollution and misses carbon stars that exhibit relatively weak molecular bands . They are mixed with non-carbon stars in the head strength map, which makes it difficult to distinguish using this method. Our proposed method can be used as an improvement of the above method, which can identify such carbon stars from the spectrum and give a characteristic attribution explanation.
@@
The organizational logic of this paper is as follows: In Section 2, we cover the source of the data, the preparation of the training set, and data processing methods. In Section 3, we introduce the proposed deep learning classification model and important model parameters, as well as the SHAP interpretability model and evaluation index respectively. In Section 4, we show how well the model performs in the validation set, how it compares to other classification models, demonstrating the effectiveness of model interpretation and key features, and the newly discovered carbon star candidates. In Section 5, We compare the results with others, analyze and interpret the spectral features. Finally, a general summary is provided in Section 6, and candidates are given in Section B.
@@ 2.1 CSTAR sample of DR3
Data Release 3 ( DR3) has been released in 2022, and its mean BP/RP spectral data was released in May of the same year, including four folders on the archive: rvs_mean_spectrum, xp_continuous_mean_spectrum, xp_sampled_mean_spectrum and xp_summary. The spectra we adopt are from xp_sampled_mean_spectrum, which has 34,468,373 BP/RP externally calibrated sampled mean spectra. All mean spectra were sampled to the same set of absolute wavelength positions, viz. 343 values from 336 to 1020 nm with a step of 2βnm, i.e. corresponding to a wavelength range of 3360 to 10200βΓ Β in steps of 20βΓ . To standardize the data scale range, we simply normalized the raw spectral data as
@@
Min-max normalization is a simple data standardization method that maps the data to a 0-1 range in equal proportion, thus changing the distribution of the raw data. It only scales the spectrum without altering its original structural representation. It also helps to speed up the convergence rate of the loss function of the deep learning model, improving the efficiency and speed of the modelβs optimal solution, which may otherwise be difficult to converge.
Gaia Collaboration etΒ al. (2023) (Hereafter referred to as C2023) released a batch of golden samples, which include the golden sample of carbon stars from βCSTARβ. Their ESP-ELS module attempted to flag these suspected carbon stars. The module is based on a random forest classifier trained on the synthetic BP and RP spectra and a sample of Galactic carbon stars (Abia etΒ al. 2020) obtained from the low-resolution spectra (; Carrasco etΒ al. 2021). In total, 386,936 targets received the βCSTARβ tag. i.e., the potential carbon star candidates. While most of these stars are M stars rather than carbon stars, only smaller fraction of the sample exhibit significant C and CN molecular bands. Most of the candidate carbon stars have - ΒΏ 2 mag and have colors consistent with M stars (Gaia Collaboration etΒ al. 2023).
As shown in Table 1, C2023 considered four molecular band heads in order to further screening reliable carbon stars.
@@
Finally, they selected 15,740 golden sample carbon stars based on the strongest CN molecular band features. Most of these golden sample carbon stars are AGB stars of which the enriched carbon was produced and then dredged up from its interior region, i.e., the cold and bright C-N type carbon stars. The calculation formula for the four molecular band head strengths is as follows:
@@
To exploit these samples, The identification results of βCSTARβ are saved in the spectraltype_esphs field of the astrophysical_parameters table. The list of βsource_idβ for golden sample of carbon stars is available in the βgold_sample_carbon_starsβ archive of the performance verification table.
@@ 2.2 Training data
Considering that the provided by DR3 may not be reliable, in Figure 1, we also plotted the and diagrams of all CSTAR_XPM in order to get a rough estimate and a basic idea of the physical properties and evolutionary state of these sources. Since a large proportion of the sources are missing and , we plotted 3245 CSTAR_XPM_G sources, 47,290 CSTAR_XPM_nonG sources, and 6763 random sample sources.
@@
It is clear to see the CSTAR_XPM_nonG and CSTAR_XPM_G are spatially intermingled in the distribution and are difficult to distinguish from each other.
@@
From the panel (a) and (c) of Figure 1, we noted that a few targets (19) have a greater than 3 and hotter than 6000 K. However, their is around 3, and their spectra are typical of AGB carbon stars (showing strong CN bands), which is quite unexpected. Although the possibility of them being extrinsic carbon stars (such as dwarf carbon stars) cannot be ruled out, C2023 mentioned that the of these targets tends to be overestimated. Therefore, it is more likely that these are due to measurement errors.
@@
In order to further ensure the purity of the carbon star sample, we visually checked all the CSTAR_XPM_G, finding that some spectra exhibit very significant contamination, which may be due to the signal-to-noise problem, or exhibit insignificant CN characteristics. These spectra would have troubled the purity of our training set. Therefore, we need to eliminate the contaminated spectra. We show the BP and RP spectra of a standard carbon star and some of the deleted problematic stars in Figure 2. After screening we selected 8176 carbon stars as the positive training sample, which showed significant CN molecular band characteristics.
@@
We then randomly selected 9,000 spectra out of 74,740 CSTAR_XPM_nonG as βnegative sampleβ, the majority of which are M-type stars (Gaia Collaboration etΒ al. 2023). First, we made sure that none of the 9,000 spectra are included in several known common carbon star lists (Solar Neighbourhood Carbon stars: Abia etΒ al. 2020; Galactic Carbon stars: Alksnis etΒ al. 2001; LMC Carbon stars: Kontizas etΒ al. 2001; Carbon stars in the SMC: Morgan & Hatzidimitriou 1995), and that none of them were labeled C* by SIMBAD or listed as βCarbonβ by LAMOSTβs pipeline. Second, here we must mention our training method. We randomly trained several models based on the above dataset. We identified false positive (FP) spectra resulting from obvious modelβs misclassification, where the true category was 0 (other stars), but the model predicted category 1 (carbon star). We iteratively added these spectra to our βnegative sampleβ dataset, finding that this data enhancement method significantly improved the modelβs generalization capability. Finally, during the visual examination of the βnegative sampleβ, we found that many spectra exhibited relatively obvious and prominent molecular bands, particularly two strong CN molecular bands. Therefore, in order to further ensure the purity of the βnegative sampleβ and ensure that the model could learn the correct features, we calculated the band head strength for all spectra according to the method provided in C2023. We screened out all the suspicious potential carbon stars under the two conditions of R and R, which are almost the weakest CN molecular band strength conditions. We obtained 2255 spectra in total, and then carefully identified the weak CN band head by eyes to exclude the interference of possible M and K type giant stars. The reason is that the absorption of M and K giants at the TiO5 (7126-7135βΓ ; Li etΒ al. 2023) position is more abundant, and may exhibit a pseudo-R molecular band strength. Through this checking process, we eliminate 885 objects that exhibit weaker CN molecular bands relative to the cold and bright golden sample of carbon stars. We showed six representative spectra in Figure 3.
@@
Note that these spectra only show weak CN molecular bands, which does not mean that they are all potential carbon stars, we just try to eliminate carbon star contamination as much as possible to obtain pure giant star sample. We also searched and excluded the problematic spectra with condition R ΒΏ 1.10 or R ΒΏ 1.30, as some of the problematic spectra may have large R or R. Then we screened the obtained spectra for three times to ensure that all spectra did not show any carbon characteristics and then added some FP spectra that apparently misclassified by the model mentioned above. Finally we kept 8556 training negative sample, which can be confidently identified as non-carbon stars. We plotted the molecular band head strength distribution of the sample in Figure 4.
@@
All the samples we used were summarised in Table 2. The Venn diagram in Figure 5 depicts the relationship between these samples.
@@
We finally got 16,732 stars as training sample, the total number of which is approximately 20.15% of the total number of CSTAR_XPM sample. We will use this training sample to verify the accuracy of our classification model and the explanatory performance of the interpretable model. We speculate that there may be some additional carbon stars in CSTAR_XPM_nonG, so we will apply the trained model to this sample.

### v2

@@ Abstract
Context. A large fraction of Asymptotic Giant Branch (AGB) stars develop carbon-rich atmospheres during their evolution. Based on their color and luminosity, these carbon stars can be easily distinguished from many other kinds of stars. However, numerous G, K, and M giants also occupy the same region as carbon stars on the HR diagram. Despite this, their spectra exhibit differences, especially in the prominent CN molecular bands.
Aims. We aim to distinguish carbon stars from other kinds of stars using βs XP spectra, while providing attributional interpretations of key features necessary for identification, and even discovering additional new spectral key features.
Methods. We propose a classification model named βGaiaNetβ, an improved one-dimensional convolutional neural network specifically designed for handling βs XP spectra. We utilized the SHAP interpretability model to determine SHAP values for each feature in a spectrum, enabling us to explain the output of the βGaiaNetβ model and provide further meaningful analysis.
Results. Compared to four traditional machine-learning methods, the βGaiaNetβ model exhibits an average classification accuracy improvement of approximately 0.3% on the validation set, with the highest accuracy reaching 100%. Utilizing the SHAP model, we present a clear spectroscopic heatmap highlighting molecular band absorption features primarily distributed around CN and CN, and summarize five key feature regions for carbon star identification. Upon applying the trained classification model to the CSTAR sample with βxp_sampled_meanβ spectra, we obtained 451 new candidate carbon stars as a by-product.
@@
Carbon stars, first recognized by Secchi (1869), exhibit an inversion of the C/O ratio (C/O ΒΏ 1). Compared to other stars, carbon stars possess carbon-enriched atmospheres. The enrichment can originate from mass transfer in binary systems or the activation of the third dredge-up (TDU) process during their AGB evolution phase (Gaia Collaboration etΒ al. 2023). The carbon stars arise in binary systems likely inherited atmospheric carbon through mass transfer from its AGB companion, which now is a white dwarf (Abia etΒ al. 2002). These carbon stars are known as βextrinsicβ carbon stars. However, the carbon enrichment of cool and luminous carbon stars (N-type) is caused by the pollution of nuclear helium fusion products transported from the inner to the outer layers during their AGB phase (Gaia Collaboration etΒ al. 2023). This process transports carbon from the interior region to the stellar surface, leaving the outer layers rich in carbon-containing molecules and dust particles. Thus, their spectra show C and CN molecular bands stronger than usual in stars cooler than 3800 K (i.e. ; Gaia Collaboration etΒ al. 2023). These carbon stars are called βintrinsicβ carbon stars.
@@
The spectra of carbon stars exhibit absorption characteristics due to carbon-containing compounds of CH, C, and CN. The presence and intensity of these spectral features provide valuable insights into the atmospheric conditions and chemical processes within these stars. Because they tend to be in a late evolution stage of stellar mass loss, carbon stars are important contributors to the interstellar medium and serve as a good references for studying a variety of physical processes affecting the end of the life of low mass stars (Gaia Collaboration etΒ al. 2023).
@@
Traditionally, identifying and classifying carbon stars relied on parametric measurements and manual checking of their spectra. Ji etΒ al. (2016) identified 894 carbon stars from LAMOST DR2 by measuring multiple line indices from the stellar spectra. Abia etΒ al. (2022) reported the identification of 2660 new carbon star candidates through 2MASS photometry, astrometry, and their location in the β2MASS diagram. Li etΒ al. (2023) distinguished carbon stars from M-type giants by selecting spectral indices satisfying the criterion []. Gaia Collaboration etΒ al. (2023) screened carbon stars by measuring the carbon molecular band head strength, resulting in a final selection of 15,740 high-quality βgolden sampleβ carbon stars. Lebzelter etΒ al. (2023) classified 546,468 carbon star candidates using a method guided by narrow-band photometry (Palmer & Wing 1982). Their SOS module examines this feature in an automated way by computing the pseudo-wavelength difference between the two highest peaks in each DR3 RP spectrum, taking the median value of the results, and storing it in the parameter of median_delta_wl_rp. They assumed all stars with median_delta_wl_rp to be C-rich stars. Li etΒ al. (2024) identified 3546 carbon stars through line indices and near-infrared colorβcolor diagrams. Through visual inspection of these spectra, they further subclassify them into CβH, CβR, CβN, and Ba stars.
@@
To date, many machine-learning methods have been used to identify carbon stars based on their spectra. Si etΒ al. (2014) applied the label propagation algorithm to search for 260 new carbon stars from SDSS (Sloan Digital Sky Survey) DR8. Si etΒ al. (2015) applied the efficient manifold ranking algorithm to search for 183 carbon stars from the LAMOST pilot survey. Li etΒ al. (2018) identified 2651 carbon stars in the spectra of more than seven million stars using an efficient machine-learning algorithm for LAMOST (Large Sky Area Multi-Object Fiber Spectroscopic Telescope) DR4. Sanders & Matsunaga (2023) investigated the use of unsupervised learning algorithms to classify the chemistry of long-period variables (Lebzelter etΒ al. 2023) from DR3βs BP/RP spectra (also called XP spectra; Carrasco etΒ al. 2021) into O-rich and C-rich groups. They also employed a supervised approach to separate O-rich and C-rich sources using broadband optical and infrared photometry. In all, they tagged a total of 23,737 C-rich classifications based on the BP/RP spectra and identified a small population of C-rich stars in the Galactic bar-bulge region.
@@
In recent years, the application of interpretable analysis techniques in astronomy has shown great potential. Qin etΒ al. (2019) used a random forest (RF) algorithm deriving rank features, then picked out 15,269 Am candidates from the early-type stars of LAMOST DR5. He etΒ al. (2022) successfully distinguished between Red Giant Branch (RGB) and Red Clump (RC) stars using the XGBoost algorithm (Chen & Guestrin 2016), and used the SHAP interpretable model (Lundberg & Lee 2017) to obtain the top features that the XGBoost selected. Shang etΒ al. (2022) used machine learning algorithms to search for class-one and class-two chemical peculiars (CP1 and CP2). Finally they presented a catalog of 6917 CP1 and 1652 CP2 new candidate sources using XGBoost followed by the visual investigation, and listed the spectral features for separating CP1 from CP2 using SHAP.
@@
Before this work, searching for carbon stars based on spectra was mainly to obtain the matching giant star spectra by random forest classifier (Breiman 2001), and then to screen out carbon stars with strong CN molecular bands by measuring the molecular band head strength (Gaia Collaboration etΒ al. 2023), most of which belong to the N-type. Although this method can easily pick out the carbon stars with obvious spectral features, it misses carbon stars that exhibit relatively weak molecular bands. They are mixed with non-carbon stars in the band head strength diagrams, which makes it difficult to distinguish them using this method. Our proposed algorithm can be used as an improvement of the above method, to identify such carbon stars from their spectra in a quantitative manner.
@@
In this work, we explored the significant potential of deep learning model enhanced with interpretability analysis algorithm (SHAP) for identifying carbon stars using their DR3 XP spectra. In Section 2, we cover the source of the data, the preparation of the training set, and data processing methods. In Section 3, we introduce the proposed deep learning classification model and important model parameters, as well as the SHAP interpretability model, and the evaluation index. In Section 4, we show how well the model performs with a validation set, how it compares to other classification models, demonstrating the effectiveness of model interpretation and key features, as well as the newly discovered carbon star candidates. In Section 5, we compare the results with others, analyze and interpret the spectral features. Finally, a general summary is provided in Section 6, and candidates are given in Section B.
@@ 2.1 CSTAR sample of DR3
Data Release 3 ( DR3) has been released in 2022, and its mean BP/RP spectral data was released in May of the same year, including four folders on the archive: rvs_mean_spectrum, xp_continuous_mean_spectrum, xp_sampled_mean_spectrum and xp_summary. The spectra we adopt are from xp_sampled_mean_spectrum, which has 34,468,373 BP/RP externally calibrated sampled mean spectra. All mean spectra were sampled to the same set of absolute wavelength positions, viz. 343 values from 336 to 1020 nm with a step of 2βnm, i.e. corresponding to a wavelength range of 3360 to 10200βΓ Β in steps of 20βΓ . To standardize the data scale range, we normalized the raw spectral data as
@@
Min-max normalization is a data standardization method that maps the data to a 0-1 range in equal proportion, thus changing the distribution of the raw data. It only scales the spectrum without altering its original structural representation. It also accelerates the convergence rate of the deep learning modelβs loss function, thereby enhancing the efficiency and speed of achieving the optimal solution, which might otherwise be hard to converge.
Gaia Collaboration etΒ al. (2023) (hereafter referred to as C2023) released a batch of βgolden samplesβ, which include the βgolden sampleβ of carbon stars from βCSTARβ. Their ESP-ELS module attempted to flag these suspected carbon stars. The module is based on a random forest classifier trained on the synthetic BP and RP spectra and a sample of Galactic carbon stars (Abia etΒ al. 2020) obtained from the low-resolution spectra (; Carrasco etΒ al. 2021). In total, 386,936 targets received the βCSTARβ tag. i.e., the potential carbon star candidates. While most of these stars are M stars rather than carbon stars, only smaller fraction of the sample exhibit significant C and CN molecular bands. Most of the candidate carbon stars have - ΒΏ 2 mag and have colors consistent with M stars (Gaia Collaboration etΒ al. 2023).
As shown in Table 1, C2023 considered four molecular band heads in order to further screen for reliable carbon stars.
@@
Finally, they selected 15,740 golden sample carbon stars based on the strongest CN molecular band features. Most of these golden sample carbon stars are AGB stars of which the enriched carbon was produced and then dredged up from its interior region, i.e., the cold and bright N-type carbon stars. The calculation formula for the four molecular band head strengths is as follows:
@@ 2.2 Training data
Considering that the provided by DR3 may not be reliable, in Figure 1, we also plotted the (not dereddened) and diagrams of all CSTAR_XPM in order to get a basic idea of the physical properties and evolutionary state of these sources. Since a large proportion of the sources are missing and , we plotted 3245 CSTAR_XPM_G sources, 47,290 CSTAR_XPM_nonG sources, and 6763 random sample sources.
@@
It is clear that the CSTAR_XPM_nonG and CSTAR_XPM_G are spatially intermingled in the distribution and are difficult to distinguish from each other.
@@
From the panel (1(a)) and (1(c)) of Figure 1, we noted that a few targets (19) have a greater than 3 and hotter than 6000 K. However, their is around 3, and their spectra are typical of AGB carbon stars (showing strong CN bands), which is quite unexpected. Although the possibility of them being extrinsic carbon stars (such as dwarf carbon stars) cannot be ruled out, C2023 mentioned that the of these targets tends to be overestimated. Therefore, it is more likely that these are due to stellar parameter determination errors.
@@
In order to further ensure the purity of the carbon star sample, we visually checked all CSTAR_XPM_G spectra. We found that some spectra exhibit significant contamination, which may be due to the signal-to-noise (S/N) problem, and some spectra exhibit insignificant CN characteristics. These spectra would have troubled the purity of our training set. Therefore, we need to eliminate these problematic spectra. We show the XP spectra of a standard carbon star and some of the deleted problematic stars in Figure 2. After screening we selected 8176 carbon stars as the positive training sample, which showed significant CN molecular band characteristics.
@@
We then randomly selected 9,000 spectra out of 74,740 CSTAR_XPM_nonG as βnegative sampleβ, the majority of which are M-type stars (Gaia Collaboration etΒ al. 2023). First, we made sure that none of the 9,000 spectra are included in several known common carbon star lists (Solar Neighbourhood carbon stars: Abia etΒ al. 2020; Galactic carbon stars: Alksnis etΒ al. 2001; carbon stars in the Large Magellanic Cloud (LMC): Kontizas etΒ al. 2001; carbon stars in the Small Magellanic Cloud (SMC): Morgan & Hatzidimitriou 1995), and that none of them were labeled C* by SIMBAD or listed as βCarbonβ by LAMOSTβs pipeline. Second, here we must mention our training method. We randomly trained several models based on the above dataset. We identified false positive (FP) spectra resulting from obvious model misclassification, where the true category was 0 (other stars), but the model predicted category 1 (carbon star). We iteratively added these spectra to our βnegative sampleβ dataset, finding that this data enhancement method significantly improved the modelβs generalization capability. Finally, during the visual examination of the βnegative sampleβ, we found that many spectra exhibited relatively obvious and prominent molecular bands, particularly the two strong CN molecular bands. Therefore, in order to further ensure the purity of the βnegative sampleβ and ensure that the model could learn the correct features, we calculated the band head strength for all spectra according to the method provided in C2023. We screened out all the suspicious potential carbon stars under the two conditions of R and R, which are almost the weakest CN molecular band strength conditions. We obtained 2255 spectra in total, and then carefully identified the weak CN band head by eye to exclude the interference of possible M and K type giant stars. The reason is that the absorption of M and K giants at the TiO5 (7126-7135βΓ ; Li etΒ al. 2023) position is more abundant, and may exhibit a pseudo-R molecular band strength. Through this checking process, we eliminate 885 objects that exhibit weaker CN molecular bands relative to the cold and bright golden sample of carbon stars. We showed six representative spectra in Figure 3.
@@
Note that these spectra only show weak CN molecular bands, which does not mean that they are all potential carbon stars, we just try to eliminate carbon star contamination as much as possible to obtain pure giant star sample. We also searched and excluded the problematic spectra with condition R ΒΏ 1.10 or R ΒΏ 1.30, as some of the problematic spectra may have large R or R. Then we screened the obtained spectra three times to ensure that all spectra did not show any carbon characteristics and then added some FP spectra that were apparently misclassified by the model mentioned above. Finally we kept 8556 training negative sample, which can be confidently identified as non-carbon stars. We plotted the molecular band head strength distribution of the sample in Figure 4.
@@
All the samples we used are summarised in Table 2. The Venn diagram in Figure 5 depicts the relationship between these samples.
@@
We finally got 16,732 stars as training sample, the total number of which is approximately 20.15% of the total number of CSTAR_XPM sample. This training sample is used to verify the accuracy of our classification model and the explanatory performance of the interpretable model. We speculate that there may be some additional carbon stars in CSTAR_XPM_nonG, so we will apply the trained model to this sample.